OpenStudy (weaver):

3. Joe solved and justified each step as shown. Identify Joe’s error and fix his mistake. Step Justification Given equation Multiplication Property of Equality Subtraction Property of Equality

OpenStudy (weaver):

@jackthegreatest .... Is this correct? 3+x/2=10 Given 3+x=20 Multiplication prop of = x=17 Subtraction prop o

OpenStudy (jackthegreatest):

do you have a screenshot

OpenStudy (jackthegreatest):

@Weaver

OpenStudy (jackthegreatest):

close, for the second step, you multiply everything by 2 so 6 + x = 20

OpenStudy (weaver):

thanks

OpenStudy (jackthegreatest):

np